OpenStudy (lisa123):

Slyanite is a mineral that contains 28.0% gold by mass. How much sylvanite would you need to dig up to obtain 83.0g of gold

OpenStudy (lisa123):

@phi

Vocaloid (vocaloid):

set up a proportion mass of gold/total mass = 28/100 that gives us 83.0/x = 28/100 solve for x

OpenStudy (anonymous):

Wouldn't it be 0.28 times 83?

Vocaloid (vocaloid):

not quite. if we wanted to know how much gold was in 83.0 grams of slyaniate, then yes, your approach would work, but the question wants to know how much sylvaniate has 83.0 g of gold

OpenStudy (lisa123):

okay thank you @vocaloid I got the answer right!

OpenStudy (phi):

***Wouldn't it be 0.28 times 83? *** no, it would be 0.28 * unknown amount of mineral = 83 g pure gold or 0.28 x= 83 solve for x this will give the same answer as using proportions / ratios